Solve the given problem releated to continuous compounding interest.
How long will it take $1,400 to triple if it is invested at an annual interest rate of 5.3% compounded continuously? Round to the nearest year.

Answer by Boreal(15235)

P=Poexp^(rt)
P/Po=3
3=exp^rt
ln3=rt
ln3/0.053=t
t=21 years